# MSCPT: Few-shot Whole Slide Image Classification with Multi-scale and Context-focused Prompt Tuning

### v1

@@ III-B1 Review of CLIP
CLIP [14] adopts a two-tower structure, including an image encoder and a text encoder. The image encoder $\mathcal{F}_{img}$ can be either a ResNet [37] or ViT [38], which is used to transform images into visual embeddings. The text encoder $\mathcal{F}_{text}$ takes a series of words as input and outputs textual embeddings. During the training process, CLIP utilizes a contrastive loss to learn a joint embedding space for the two modalities. During inference, we assume $\boldsymbol{x}$ is the visual embedding, and $\left\{\boldsymbol{w_{i}}\right\}_{i=1}^{K}$ is a series of textual embeddings generated by $\mathcal{F}_{text}$. Each $\boldsymbol{w_{i}}$ corresponds to prompt (such as “an image of {class name}) embedding for a specific image category. Therefore, the predicted probabilities can be obtained by calculating the cosine similarity between $\boldsymbol{x}$ and $\boldsymbol{w_{i}}$:
$p(y=i\mid\boldsymbol{x})=\frac{exp(cos(\boldsymbol{x},\boldsymbol{w_{i}})/\tau)}{{\textstyle\sum_{j=1}^{K}exp(cos(\boldsymbol{x},\boldsymbol{w_{j}})/\tau)}},$ (2)
@@ III-B2 Patch Selection
Due to the high resolution of WSIs, dividing them into non-overlapping patches will result in a large number of patches. However, research has shown that only a few patches contain crucial information [31]. By preliminarily identifying patches closely linked to cancer analysis, we can notably diminish the computational resources demanded by visual prompt tuning. The powerful zero-shot ability of the VLMs allows for the initial screening of cancer-related patches.
@@
Specifically, we utilize $\mathcal{F}_{img}$ to extract visual embeddings from patches while leveraging $\mathcal{F}_{text}$ to extract textual embeddings from the category prompts. Following this, the similarities between patches and prompts are computed. Then, the top $n$ patches with the highest similarity scores are selected for each category. To enhance the robustness of patch selection, we generated 50 sets of manual category templates and averaged their embeddings following  [21]. For a WSI $X_{i}$, we choose patches $\left\{x^{l}_{i,j},j=1,2,...,n_{l}\right\}$ at low magnification. Due to our unique architecture, we solely perform patch selection and visual prompt tuning at low magnification.
III-C Multi-scale Visual Descriptions Construction
@@
In this part, we aim to generate pathological visual descriptions as pathological language prior knowledge to guide the hierarchical prompt tuning and instances aggregation. To reduce manual workload, large language models (LLMs) are employed to generate descriptions related to different diseases. That is, we enter the question “We are studying Cancer Category. Please list $\boldsymbol{C^{l}}$ visual descriptions at 5×5$\times$5 × magnification and $\boldsymbol{C^{h}}$ visual descriptions at 20×20$\times$20 × magnification observed in H&E-stained histological images of Cancer Sub-category.” into the LLM. And then we can get the multi-scale visual description sets $T^{low}=\{T^{low}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{l}\}$ and $T^{high}=\{T^{high}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{h}\}$. $K$ represents the number of WSI categories, and $C^{l}$ and $C^{h}$ denote the counts of low-level and high-level descriptions, respectively.
III-D Multi-scale Hierarchical Prompt Tuning
@@
Inspired by MetaPrompt [26], a unique dual-path asymmetric framework is employed for multimodal hierarchical prompt tuning, as shown in the left of Fig. 2. Freezing two of all encoders helps reduce the trainable parameters and alleviates overfitting in few-shot scenarios. Compared to previous works where their encoders process the same inputs, our method adopts a unique strategy: the prompted and frozen encoders take entirely different inputs. Considering the immense size of WSIs and the substantial computational and storage resource requirements for visual prompt tuning, we only conducted visual prompt tuning at the low level.
@@
Rather than modifying the visual prompts tuning method from Metaprompt, our emphasis is placed on the text modality. Specifically, the low-level pathological visual descriptions $T^{low}$ are sent into the frozen low-level text encoder $\mathcal{F}^{low}_{text}$, while the high-level pathological visual descriptions $T^{high}$ are sent into the prompted hierarchical high-level text encoder $\mathcal{F}^{high}_{text}$. Simultaneously, patches are also fed into the corresponding encoders. We wish to integrate different information contained in $T^{low}$ and $T^{high}$, which can help improve the multi-scale information processing capability of MSCPT. To achieve this purpose, we propose Multi-scale Hierarchical Prompt Turning (MHPT) module. The core component of MHPT, prompted hierarchical high-level text encoder, has been drawn in Fig. 3.
@@
III-D1 Multi-scale Prompts Construction
@@
$p_{low,i}^{l}=g\left(d_{low}^{l}\right),$ (3)
where $d_{low}^{l}$ is the last token of $T^{low}_{k,i}$ at the $l$-th layer, generator $g$ is a basic multilayer perceptron to align vectors of different scales into a common embedding space. Then, these tokens get concatenated to obtain low-level text prompts $p_{low}^{l}$.
III-D2 Hierarchical Prompt Tuning
@@
$z^{high}_{k,c}=Textproj(E^{L}).$ (5)
III-E Image-text Similarity-based Graph Prompt Tuning
Some studies have shown that the interactions between different areas of WSI and their structural information play a crucial role in cancer analysis [28]. However, the current prompt tuning methods are unable to capture this information. To address this, we propose Image-text Similarity-based Graph Prompt Tuning (ISGPT) module. More specifically, we deviate from conventional methods that utilize patch coordinates or patch feature similarity in constructing graph neural networks (GNNs) [29, 30]. Our innovative approach involves utilizing the similarity between patches and pathological visual descriptions as the foundation for developing GNNs. We treat the patches as nodes and aim to construct the adjacency matrix $\boldsymbol{A}$ by calculating the semantic similarity $\boldsymbol{S}$ between the patch embeddings and description embeddings. Specifically, after patches and descriptions have passed through the encoders from Section III-D, patch embeddings $\boldsymbol{P}\in\mathbb{R}^{M\times d}$ and description embeddings $\boldsymbol{Z}\in\mathbb{R}^{KC\times d}$ are obtained, respectively. The formula for semantic similarity $\boldsymbol{S}\in\mathbb{R}^{M\times KC}$ is:
$s_{i,j}=\frac{exp(cos(\boldsymbol{P_{i}},\boldsymbol{Z_{j}})/\tau)}{{\textstyle\sum_{m=1}^{K\times C}exp(cos(\boldsymbol{P_{i}},\boldsymbol{Z_{m}})/\tau)}},$ (6)
@@ III-E Image-text Similarity-based Graph Prompt Tuning
$a_{i,j}=\frac{exp(cos(\boldsymbol{S_{i}},\boldsymbol{S_{j}})/\tau)}{{\textstyle\sum_{m=1}^{M}exp(cos(\boldsymbol{S_{i}},\boldsymbol{S_{m}})/\tau)}},$ (7)
where $\boldsymbol{S_{i}}\in\mathbb{R}^{KC}$ represents the semantic similarity between $i$-th patch embeddings and all description embeddings(i.e., the $i$-th row of $\boldsymbol{S}$). We avoid constructing $\boldsymbol{A}$ based on patch coordinates or patch feature similarity, as this approach might overlook fewer but significant patches when focusing only on Euclidean distance or patch feature similarity. Subsequent experimental results have demonstrated the superior performance of our method for constructing $\boldsymbol{A}$. We choose Graph Convolutional Network (GCN) [39] as the graph learning model. The definition of the GCN operation in the $l$-th GCN layer is as follows:
$\mathcal{F}_{GCN}\left(\boldsymbol{A},\boldsymbol{H}^{\left(l\right)}\right)$ $=\sigma(\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\tilde{\boldsymbol{A}}\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\boldsymbol{H^{\left(l\right)}}\boldsymbol{W}^{\left(l\right)}).$ (8)
@@
$\tilde{\boldsymbol{P}}^{high}$ $=\mathcal{F}^{high}_{GCN}(\boldsymbol{A}^{high},\boldsymbol{P}^{high}),$ (9)
@@
III-F Non-Parametric Cross-Guided Pooling
Impressed by the powerful zero-shot capability of pre-trained VLMs, the possibility of employing a similar non-parametric approach for instance aggregation was pondered. We propose Non-Parametric Cross-Guided Pooling (NPCGP) to aggregate instance into bag features. In NPCGP, we compute semantic similarities between the patch embeddings $\tilde{\textbf{P}}$ post graph tuning and pathological visual description embeddings $\boldsymbol{Z}$ at both the same and across scales. The reason for calculating similarities both within the same and across scales is our concern that the pathological visual descriptions provided by LLM may contain scale-related inaccuracies. Hence, this procedure serves to bolster the robustness of feature aggregation strategies. Lastly, the bag-level unnormalized probability distribution $Logits$ is obtained through the topK max-pooling operator $h_{topK}$:
@@ III-F Non-Parametric Cross-Guided Pooling
Following previous work [26], we use cross-entropy loss to optimize the three distributions $Logits^{overall}$, $Logits^{high}$, and $Logits^{low}$, but only $Logits^{overall}$ was used during model inference.
TABLE I: Cancer sub-typing results on TCGA-NSCLC, TCGA-BRCA, and TCGA-RCC. The highest performance is in bold, and the second-best performance is underlined. We provided mean and standard deviation results under five random seeds.

### v2

@@ 3.2.1 Review of CLIP
CLIP [12] adopts a two-tower structure, including an image encoder and a text encoder. The image encoder $\mathcal{F}_{img}$ can be either a ResNet [37] or a ViT [38], which is used to transform images into visual embeddings. The text encoder $\mathcal{F}_{text}$ takes a series of words as input and outputs textual embeddings. During the training process, CLIP utilizes a contrastive loss to learn a joint embedding space for the two modalities. During inference, we assume $\boldsymbol{x}$ is the visual embedding, and $\left\{\boldsymbol{w}_{i}\right\}_{i=1}^{K}$ is a series of textual embeddings generated by $\mathcal{F}_{text}$. Each $\boldsymbol{w}_{i}$ corresponds to a prompt (such as “an image of {class name}) embedding for a specific image category. Therefore, the predicted probabilities can be obtained by calculating the cosine similarity between $\boldsymbol{x}$ and $\boldsymbol{w}_{i}$:
$\small p(y=i\mid\boldsymbol{x})=\frac{exp(cos(\boldsymbol{x},\boldsymbol{w}_{i})/\tau)}{{\textstyle\sum_{j=1}^{K}exp(cos(\boldsymbol{x},\boldsymbol{w}_{j})/\tau)}},$ (2)
@@ 3.2.2 Patch Selection
Some studies have shown that simultaneously performing prompt tuning on the visual encoder while tuning the text encoder significantly improves the performance of VLMs across various downstream tasks [14, 23]. Due to the high resolution of WSIs, dividing them into non-overlapping patches will result in a large number of patches. Performing visual prompt tuning on all patches is computationally impractical due to the significant resource requirements. Fortunately, research has shown that only a few patches contain crucial information [28]. By leveraging the robust zero-shot capabilities of VLMs to preliminarily identify patches most relevant to cancer analysis, the computational cost of visual prompt tuning can be significantly reduced.
@@
Specifically, we utilize $\mathcal{F}_{img}$ to extract visual embeddings from patches while leveraging $\mathcal{F}_{text}$ to extract textual embeddings from the category prompts. For each class, a category prompt consists of two components: the class name (e.g., “squamous cell carcinoma”) and the template (e.g., “a photomicrograph of {class name}), which together form the final prompt (e.g., “a photomicrograph of squamous cell carcinoma”). To enhance the robustness of patch selection, we randomly generated 50 sets of prompts using various category names and templates. Following the method in [18], we averaged the embeddings of these prompts. Subsequently, the similarities between patches and prompts are computed. Then, the top $n$ patches with the highest similarity scores are selected for each category. For a WSI $X_{i}$, we choose patches $\left\{x^{l}_{i,j},j=1,2,...,n_{l}\right\}$ at low magnification. Due to our unique architecture, we solely perform patch selection and visual prompt tuning at low magnification.
3.3 Multi-scale Visual Descriptions Construction
@@
In this part, we aim to generate pathological visual descriptions as pathological language prior knowledge to guide the hierarchical prompt tuning and instances aggregation. To reduce manual workload, large language models (LLMs) are employed to generate descriptions related to different diseases. That is, we enter the question “We are studying Cancer Category. Please list $\textbf{C}^{l}$ visual descriptions at 5×5$\times$5 × magnification and $\textbf{C}^{h}$ visual descriptions at 20×20$\times$20 × magnification observed in H&E-stained histological images of Cancer Sub-category.” into the LLM. And then we can get the multi-scale visual description sets $T^{low}=\{T^{low}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{l}\}$ and $T^{high}=\{T^{high}_{k,c}\mid 0\leq k\leq K,0\leq c\leq C^{h}\}$. $K$ represents the number of WSI categories, and $C^{l}$ and $C^{h}$ denote the counts of low-level and high-level descriptions, respectively.
3.4 Multi-scale Hierarchical Prompt Tuning
@@
Inspired by MetaPrompt [23], a unique dual-path asymmetric framework is employed for multimodal hierarchical prompt tuning, as shown on the left of Fig. 2. Freezing two of the four encoders helps reduce the trainable parameters and alleviates overfitting in few-shot scenarios. Compared to previous works where their encoders process the same inputs, our method adopts a unique strategy: the prompted and frozen encoders take entirely different inputs. Considering the immense size of WSIs and the substantial computational and storage resource requirements for visual prompt tuning, we only conducted visual prompt tuning at the low level.
@@
Rather than modifying the visual prompts tuning method, our emphasis is placed on the text modality. Specifically, the low-level pathological visual descriptions $T^{low}$ are sent into the frozen low-level text encoder $\mathcal{F}^{low}_{text}$, while the high-level pathological visual descriptions $T^{high}$ are sent into the prompted hierarchical high-level text encoder $\mathcal{F}^{high}_{text}$. Although the low-level text encoder remains frozen, we propose a novel cross-level prompt generator that projects embeddings from the low-level encoder into the prompted hierarchical high-level text encoder during training. This design preserves the information encapsulated in low-level visual descriptions. Simultaneously, patches are also fed into the corresponding encoders. We wish to integrate different information contained in $T^{low}$ and $T^{high}$, which can enhance the multi-scale information processing capability of MSCPT. To achieve this purpose, we propose Multi-scale Hierarchical Prompt Tuning (MHPT) module. The core component of MHPT, the prompted hierarchical high-level text encoder, is illustrated in Fig. 3.
@@
3.4.1 Multi-scale Prompts Construction
@@
$\small p_{low,i}^{l}=g\left(d_{low}^{l}\right),$ (3)
where $d_{low}^{l}$ is the last token of $T^{low}_{k,i}$ at the $l$-th layer, and generator $g$ is a basic multilayer perceptron to align vectors of different scales into a common embedding space. Then, these tokens get concatenated to obtain low-level text prompts $p_{low}^{l}$.
3.4.2 Hierarchical Prompt Tuning
@@
$\small z^{high}_{k,c}=Textproj(E^{L}).$ (5)
3.5 Image-text Similarity-based Graph Prompt Tuning
Some studies have shown that the interactions between different areas of WSI and their structural information play a crucial role in cancer analysis [25]. However, the current prompt tuning methods are unable to capture this information. To address this, we propose Image-text Similarity-based Graph Prompt Tuning (ISGPT) module. More specifically, we deviate from conventional methods that utilize patch coordinates or patch feature similarity in constructing graph neural networks (GNNs) [26, 27]. Our innovative approach involves utilizing the similarity between patches and pathological visual descriptions as the foundation for developing GNNs. We treat the patches as nodes and aim to construct the adjacency matrix $\boldsymbol{A}$ by calculating the semantic similarity $\boldsymbol{S}$ between the patch embeddings and description embeddings. Specifically, after patches and descriptions have passed through the encoders from Section 3.4, patch embeddings $\boldsymbol{P}\in\mathbb{R}^{M\times d}$ and description embeddings $\boldsymbol{Z}\in\mathbb{R}^{KC\times d}$ are obtained, respectively. The formula for semantic similarity $\boldsymbol{S}\in\mathbb{R}^{M\times KC}$ is:
$\small s_{i,j}=\frac{exp(cos(\boldsymbol{P}_{i},\boldsymbol{Z}_{j})/\tau)}{{\textstyle\sum_{m=1}^{K\times C}exp(cos(\boldsymbol{P}_{i},\boldsymbol{Z}_{m})/\tau)}},$ (6)
@@ 3.5 Image-text Similarity-based Graph Prompt Tuning
$\small a_{i,j}=\frac{exp(cos(\boldsymbol{S}_{i},\boldsymbol{S}_{j})/\tau)}{{\textstyle\sum_{m=1}^{M}exp(cos(\boldsymbol{S}_{i},\boldsymbol{S}_{m})/\tau)}},$ (7)
where $\boldsymbol{S}_{i}\in\mathbb{R}^{KC}$ represents the semantic similarity between $i$-th patch embeddings and all description embeddings (i.e., the $i$-th row of $\boldsymbol{S}$). We avoid constructing $\boldsymbol{A}$ based on patch coordinates or patch feature similarity, as this approach might overlook sparse but significant patches when focusing solely on Euclidean distance or feature similarity. Subsequent experimental results demonstrate the superior performance of our method for constructing $\boldsymbol{A}$. We choose a Graph Convolutional Network (GCN) [39] as the graph learning model. The GCN operation in the $l$-th layer is defined as:
$\small\mathcal{F}_{GCN}\left(\boldsymbol{A},\boldsymbol{H}^{\left(l\right)}\right)$ $=\sigma(\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\tilde{\boldsymbol{A}}\tilde{\boldsymbol{D}}^{-\frac{1}{2}}\boldsymbol{H}^{\left(l\right)}\boldsymbol{W}^{\left(l\right)}).$ (8)
@@
$\small\tilde{\boldsymbol{P}}^{high}$ $=\mathcal{F}^{high}_{GCN}(\boldsymbol{A}^{high},\boldsymbol{P}^{high}),$ (9)
@@
3.6 Non-Parametric Cross-Guided Pooling
Motivated by the powerful zero-shot capability of pre-trained VLMs, the possibility of employing a similar non-parametric approach for instance aggregation was pondered. We propose Non-Parametric Cross-Guided Pooling (NPCGP) to aggregate instances into bag features. In NPCGP, we compute semantic similarities between the patch embeddings $\tilde{\textbf{P}}$ post graph tuning and pathological visual description embeddings $\boldsymbol{Z}$ at both the same and across scales. The reason for calculating similarities both within the same and across scales is our concern that the pathological visual descriptions provided by LLM may contain scale-related inaccuracies. Hence, this procedure serves to bolster the robustness of feature aggregation strategies. Lastly, the bag-level unnormalized probability distribution $Logits$ is obtained through the Top-K max-pooling operator $h_{topK}$:
@@ 3.6 Non-Parametric Cross-Guided Pooling
Following [23], we use cross-entropy loss to optimize the three distributions $Logits^{overall}$, $Logits^{high}$, and $Logits^{low}$, but only $Logits^{overall}$ was used during model inference.
Table 1: Results on five datasets, three downstream tasks and three VLMs under the 16-shot setting. The highest performance is in bold, the second-best underlined, with mean and standard deviation reported across five random seeds.